OpenStudy (anonymous):

g(x)=5x+9/2x^2+93 Find domain of g

OpenStudy (hlambach):

The domain is all the values x can be. The only thing that cannot happen to this function is the denominator (numbers bellow/to the right of the division sign) equaling 0, because then it would be undefined. So you need to find the value(s) of x for when the denominator equals 0.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

?

OpenStudy (hlambach):

2x^2+93=0 Solve for x. Hint: if possible factor it out.

OpenStudy (hlambach):

Sorry, my internet is slow.

OpenStudy (hlambach):

The answers you get means the domain is all numbers except for (the answers you get for x). Get it?

OpenStudy (anonymous):

okay that makes sense. how do i factor that? I'm bad at factoring unless its ax^2+bx+c

OpenStudy (hlambach):

Yeah, that's way too complicated to factor this out. Just do it the regular way.

OpenStudy (anonymous):

regular way?

OpenStudy (hlambach):

Just solve for x, haha. :)

OpenStudy (hlambach):

Y'know, this whole thing 2x^2+93=0 -93 -93 ------------ 2x^2=-93 Now do the rest.

OpenStudy (anonymous):

but then I get a decimal ha. I can't use decimals. I have no idea how to square root a fraction??

OpenStudy (hlambach):

Then just leave it as a square root. remember it's plus or minus square root.

OpenStudy (hlambach):

2x^2=-93 /2 /2 --------- x^2 = -46.5 sqrt sqrt \[x=\pm \sqrt{-46.5}\] \[x=\pm i \sqrt{46.5}\]

OpenStudy (hlambach):

Get it? Okay, it is late and I'm tired. Sorry I have to abruptly leave... If you don't get it, you can tag someone else. It's just I am seriously so tired. Good nigh! :)